Find the coefficient of x12 in the expansion of each of the following.

(a) (1 − x)−7

(b)( 1 − 2x + 3x2) (1 − x)−2
Answer #1

we have (1-x)n = Ithat henti a Inenty (042) x² +.... 1:2 + 1s 2-3 n(nt) (0+2).on(ntr-1). p ! & -: is Coeft. of x2 in Clan,-zo By expansion of Cragh, we have (1-0)-2 = 1+ 2x + 3rr tualto-- - L (1-2n +32° ) (-0) To il 12 2a + 13x - - t u 1tzatan = (1-